OpenStudy (anonymous):

HOW DO YOU FIND THE DOMAIN: find the domain of f (x) : f(x)= 4/ x+3

OpenStudy (anonymous):

if f(x)=4/(x+3)...then x+3 is not =0. or x is not = -3. so domain [real numbers]-{3}. and if f(x)= 4/x +3 , domain [all real numbers]-{0}

OpenStudy (anonymous):

coudl you explain this to me, how do you know that.

OpenStudy (anonymous):

for any rational number p/q, q \[\neq\]0. so denominator of function is not =0